Mary Jarvis is a single individual who is working on filing her tax return for the previous year. She has assembled the following relevant information: She received $125,000 in salary. She received $15,500 of dividend income. She received $8,300 of interest income on Home Depot bonds. She received $23,500 from the sale of Disney stock that was purchased 2 years prior to the sale at a cost of $6,500. She received $9,500 from the sale of Google stock that was purchased 6 months prior to the sale at a cost of $4,200. Mary receives one exemption ($4,000), and she has allowable itemized deductions of $7,500. These amounts will be deducted from her gross income to determine her taxable income. Assume that her tax rates are based on Table 3.5. What is Mary's federal tax liability? Round to the nearest cent. Do not round intermediate calculation. What is the marginal tax rate? Round one decimal place. What is the average tax rate? Round to two decimal places
Answer #1

Calculation of taxable income

Income from salary = $125,000

Dividend received = $15,500

Long term capital gain = $23,500 - $6,500 = $17,000

Short Term capital gain = $9,500 - $4,200 = $5,300

Total Income = $125,000 + $15,500 + $8,300 + $17,000 + $5,300 = $171,100

Mary receives one exemption ($4,000), and she has allowable itemized deductions of $7,500.

Taxable Income = Gross Income - Exemption - Itemized Deductions = $171,000 - $4,000 - $7,500 = $159,600

Tax on Dividend Income = $15,500 *15% = $2,325

Tax on Long term Capital Gain = $17,000 * 15% = $2,550

Other Income = $159,600 - $15,500 - $17,000 = $127,100

Tax on Other Income as per 2019 federal tax brackets = $14,382.50 + 24%($127,100 - $84,200) = $24,678.5

Federal Tax Liability = Tax on Dividend Income + Tax on Long term Capital Gain + Tax on Other Income = $29,553.5

Since, the income falls under the tax bracket of 24%, Marginal Tax rate = 24%

Average Tax rate = Tax Liabilty/ Taxable Income = $29,553.5/$159,600 = 18.52%

Federal Tax Liability = $29,553.5

Marginal Tax rate = 24%

Average Tax rate = 18.52%